Christina purchased a new car and paid $12,880 after a price reduction of 8% was applied to the original price. What was the original price of the vehicle before the discount was applied? To earn credit, you MUST construct an algebraic equation with a variable and solve using algebra steps.
Answer by Theo(13342)

let x = original price of the car.
x - .08*x = 12880
.92*x = 12880
x = 12880/.92
x = 14000
-----
original price of the car is 14000.
8% of 14000 = 1120
14000 - 1120 = 12880
-----